a billiard ball is given a sharp impulse at a height h above its centre.the ball moves with a velociy v.prove tha the ball eventually acquires a velocity 1.28v if h=0.8r,r being radius of sphere


 

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

Let linear impulse=J,angular impulse=Jh

let final velocity be vf and final angular velocity be wf and initial values be v and w

vi=v, J=v+vf,wi=0, Jh=Iw, vf=r(wf), if you solve the system of equations, you will get the answer.